Your are the financial manager of Banana company, and the company want invest f this project is about 100 millions RMB which include 90 million RMB (nvestment for fixed asset and 10 millions for working capital. The tax law permIt Banana depreciate all tired asset in 5 years in straight-line method, and the marginal rate is 40% . According to your experience, you forecast the sales of new project in future 5 years as described in table 1, and from the 6th year you can get a perpetual cash flow as in 5th year, the cost of sales is about 60% of total sales and the working capital is about 10% of the sales. Now you are thinking about how to finance for this project, a new project, the initial inves you know the proportion of debt of Banana company is 40%, the interest rate is 5% and the expected return rate of equity is 15%. But you want borrow more for new project, your ideal proportion of debt of the project is 60%, and you know the interest rate will not change for good credit of Banana company. Asa financial manager, you should answer some question as follows: (10%) Question A: Calculate the cash flow of the project. Question B: Calculate the capital cost( or WACC) of the project. Question C: Calculate the NPV of the project. Sales 50100 120 10130 Cost
Answer #1

Year Sales Cost of sales EBDITA Depreciation EBIT Interest EBT Tax Profit after tax Depreciation Cash flow before working cap23 24 25 WACC Capital Debt Equity 60.00% 40.00% 3.00% 15.00% 1.80% 6.00% 7.80% 27 28 29 30 31 32 NPV 2 29.2 4 33.6 Project -15 A -F6+1 -G6+1 -H6+1 120 =1740.6 -17-18 ะเ00/5 -19-110 =10060965% -16+1 110 -J7 0.6 Year Sales Cost of sales EBDITA Depreciation EBIT Interest EBT Tax Profit after tax 130 -K7 0.6 -G7 0.6 -G7-G8 -H7-H8 ะเ00/5 -H9-H10 17-18 -100/5 K9-K10 G9-G10 -100 60965% -G11-G12 G13 0.4 -G13-G14 G10 SUM(G15:G16) -G7 10%-10 -G17-G18 -19-110 -100 60965% -J11-J12 12 100本60965% 3H11-H12 -H130.4 -H13-H14 -H10 #5UM(H15:H16) 4H7-G7) * 10% -H17-H18 1006096594 -K11-K12 K13 0.4 -K13-K14 -K10 -SUM(K15:K1 -(K7-17)10% -K17-K18 :113 *0.4 -113-114 =110 -SUM(115116) 130.4 -J13-J14 14 16 17 Cash flow before working capital Working capital cash flow Cash flovw (17-H7)10% ::117-ו18 #5UM(J15016) (17-17)*10% J17-J18 2123 B 24 25 26 27 28 29 30 C 31 32 WACC Capital Debt Equity 0.6 0.4 0.03 0.15 -G26 F26 -G27 F27 :SUM (H26:H27) -F30+1 -G19 NPV